OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 6x2 – 36x + 42 = 0

OpenStudy (anonymous):

6x^2 -36x + 42 = 0 6 (x^2 - 6x + 7) = 0 x^2 - 6x + 7 =0 x^2 - 6x = -7 x^2 - 6x + 9 = -7 + 9 (x -3)^2 = 2

jhonyy9 (jhonyy9):

moneybird you have missed the first term of 6

jhonyy9 (jhonyy9):

for you can getting this equation like resulte

OpenStudy (anonymous):



OpenStudy (anonymous):

oh i tried doing it and i got zero is that right

OpenStudy (anonymous):

(x-3)^2=2

OpenStudy (anonymous):

This is the form that you requested.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops.... I mean (x-3)^2=2